OpenStudy (anonymous):

Which of the following expressions represents the volume of a cube that is 2x on an edge? A.) 6x B.) 6x(3 squared) C.) 8x(3 squared) D.) 24x(squared)

OpenStudy (anonymous):

@Ddaddysgirl14

OpenStudy (welshfella):

volume of a cube = the edge length cubed if edge length = x then volume = x^3

OpenStudy (anonymous):

so 6x!?

OpenStudy (welshfella):

so you need to find (2x) cubed

OpenStudy (welshfella):

no 2x cubed = 2x * 2x * 2x

OpenStudy (anonymous):

=8x

OpenStudy (welshfella):

no x*x*x = x^3

OpenStudy (anonymous):

=24x

OpenStudy (welshfella):

your answered don't really make sense correct is 8x^3